# Redefining Fitness: Evolution as a Dynamic Learning Process

### v1

@@ Abstract
Evolution is the process of optimal adaptation of biological populations to their living environments. This is expressed via the concept of fitness, defined as relative reproductive success. However, it has been pointed out that this definition is incomplete and logically circular. To address this issue, several authors have called for new ways to specify fitness explicitly in terms of the relationship between phenotypes and their environment. Here, we show that fitness, defined as the likelihood function that follows from mapping population dynamics to Bayesian learning, provides a general solution to this problem. We show how probabilistic models of fitness can easily be constructed in this way, and how their averages acquire meaning as information. We also show how this approach leads to powerful tools to analyze challenging problems of evolution in variable environments, game theory, and selection in group-structured populations. The approach is general and creates an explicit bridge between population dynamics, statistical learning theory, and emerging models of artificial intelligence.
@@ Introduction
Evolutionary modeling and theory have advanced rapidly over the past century incorporating the theory of replicators [1, 2, 3, 4], sources of intragenomic variation [5, 6, 7], environmental mechanisms of genetic variation [8, 9], and socio-ecological interactions via game theory [10] and the Price equation[11].
@@
All of these approaches rely on the concept of fitness. Typically, fitness is defined as reproductive success, estimated by ratios of offspring to parents [1, 12]. Higher relative fitness of a type results in higher prevalence in the population over time. Most studies of evolutionary dynamics adopt this framework to identify better-adapted phenotypes in natural settings in theoretical models [13, 14, 15] and laboratory studies of gene expression  [16].
@@
However, it was pointed out that this definition of fitness is incomplete and logically circular. This is because it does not explicitly refer to environmental adaptation [17, 18]. A related issue deals with the nature of adaptive capacity, which must be a persistent property of phenotypes in variable environments and cannot depend on single events [16, 19]. Thus, any solution to this problem must include how fitness is averaged over time and environmental states. This issue also surfaces when modeling strategic interactions, such as in evolutionary game theory [20, 21].
@@
Finally, and related to temporal and population averaging, is the question of how fitness combines across scales. It is well known in replicator models that contributions to fitness multiply [12]. This raises questions about the approximations to fitness as a linear function of traits in both modeling and empirical analysis [22, 23]. The linearity of contributions to fitness is often justified in situations of weak selection. This is commonly assumed in the analysis of kin and group selection [22, 24] and in gene expression experiments [25]. Because many results and methods are based on the linearity of fitness, it becomes important to place this assumption on a stronger basis and to consider circumstances when selection is not weak [26, 27, 28].
@@
Other problems arise from explicit environmental or social selection, for example, in the context of evolutionary game theory. An illustration is the difficulty of deriving definite predictions from strategic interaction games for the problem of cooperation or altruism [29, 24, 30]. A related problem is the proliferation of Hamilton-type rules for similar models [31]. These complications have led to the extensive use of computer simulations with different fitness definitions [32, 33], network and group structures [34, 35], and strategy spaces [36, 37, 38, 39]. This body of work has emphasized repeated interactions [40, 41] and different types of reciprocity [31, 29] as fundamental conditions for cooperation to emerge and the need for group formation and intergroup competition [42, 22, 43]. However, these findings remain contested [24], in part because of the difficulty of analyzing the impact of assumptions and modeling choices.
@@
Here, we show that these conceptual and practical problems share a common root in the theoretical foundations of fitness. We demonstrate how they can be addressed by re-defining fitness as an optimal statistical adaptation to a probabilistic environment. Population dynamics under selection are then identified with iterated Bayesian updates of type frequencies. This identification has been recently proposed by several authors [44, 45, 46, 47], who noted its mathematical consistency and some of its conceptual implications. We take these ideas further by showing how they lead to the interpretation of fitness as a Bayesian likelihood, which is the optimal predictor of environmental states. We also show how fitness hence constructed allows us to build models of phenotype-environment selection in simple and clear ways in a variety of complex problems. This includes the effects of probabilistic environments in task selection, evolutionary dyadic games, and selection in group-structured populations.
@@
Phase spaces for these problems predict evolutionary transitions, expressed in the language of probability and information. Perhaps most satisfying, this approach shows how selection in evolutionary dynamics is a form of optimal learning theory in populations and how fitness maximization corresponds to the fundamental principle of maximizing information between population structure and the statistics of ecologically relevant environments.
@@ Results
Population replicators and their growth dynamics
We define a structured population evolving under replicator dynamics. The population is made up of different types, $s$, such that the total population size $N=\sum_{s}N_{s}$, where $N_{s}$ is the population of each type at a given time. The temporal evolution follows the standard replicator equation for a haploid population in discrete time $\tau$,
$N_{s}(t)=w_{s}N_{s}(t-\tau),\quad{\rm so\ that}\qquad N(t)=\sum_{s}N_{s}(t)=\sum_{s}w_{s}N_{s}(t-\tau)=wN(t-\tau),$ (1)
where $w_{s}\geq 0$ is the type’s fitness. The type $s$ will grow if $w_{s}>1$ and decrease for $w_{s}<1$. This leads to the growth equation for the total population with average fitness $w=\sum_{s}w_{s}f_{s}$. Here, $f_{s}=\frac{N_{s}}{N}$ is the fraction of the population of type $s$. The total population will grow if $w>1$. The growth rate, $\gamma_{s}$, is related to fitness by exponentiation $w_{s}=e^{\gamma_{s}\tau}\simeq 1+\gamma_{s}\tau$, and, correspondingly $\gamma_{s}=\frac{1}{\tau}\ln w_{s}\simeq\frac{w_{s}-1}{\tau}$. These relationships apply in continuous time while the approximations are first-order series expansions in $\tau$, Eq. 1. In the following, we set $\tau=1$ for simplicity.
@@
Fitness refers to optimal adaptation to an external environment, $e$. However, Eq. 1 and the standard set up of population genetics [1] make no explicit reference to $e$ We now show how these perspectives can be reconciled .
@@
$f_{s}(t)=\frac{N_{s}(t)}{N(t)}=\frac{w_{s}}{w}\frac{N_{s}(t-1)}{N(t-1}=\frac{w_{s}}{w}f_{s}(t-1)\qquad\rightarrow\qquad p(H|D)=\frac{p(D|H)}{p(D)}p(H).$ (2)
A number of authors have recently noted that the replicator equation under pure selection maps exactly to Bayes relation [47, 46], written on the right side of Eq. 2. Bayes theorem gives the exact update of an initial probability that a hypothesis $h\in H$ is true, $p(h)$ (the prior), given new evidence $D$. This is written as the updated (posterior) probability $p(H|D)$. Because Bayes’ theorem is an identity that follows from the rules of probability, its procedure is the optimal way to incorporate new evidence to update the initial probability [48]. As $p(H|D)$ is more constrained than $p(H)$ the data effectively selects specific hypotheses as more likely in a process where information was learned about the hypothesis space. The key quantity in this procedure is $p(D|H)$, known as the likelihood. Although it looks like another probability, the likelihood evaluated over the data and the hypotheses is the engine of selection. It is interpreted in Bayesian estimation as a probabilistic model predicting the data given each hypothesis. The remaining factor is a normalization and does not introduce any new quantity because $p(D)=\sum_{H}p(D|H)p(H)$.
@@
This identification redefines fitness as an explicit model of environmental adaptation without additional assumptions. To see this, let us map the data, $D$, to environmental states, $D\rightarrow e$, and hypotheses to population types, $H\rightarrow s$. We then identify $f_{s}(t)=f(s|e)$, where $e$ is the state of the environment at time $t$.
$f(s|e)=\frac{p(e|s)}{{\hat{p}}(e)}f(s),\rightarrow w_{s}(e)=A(e)p(e|s),\quad w(e)=\sum_{s}w_{s}(e)f(s)=A(e)\sum_{s}p(e|s)f(s)=A(e){\hat{p}}(e).$ (3)
This leads to the identification of fitness with the Bayesian likelihood, $w_{s}(e)=A(e)p(e|s)$, up to a multiplicative constant $A(e)$. The factor of $A(e)$, is necessary because fitness, defined in Eq. 1, $w_{s}=N_{s}(t)/N_{s}(t-1)>1$, when there is growth. The factor $A$ disappears in relative fitness, but contributes to absolute fitness and to the growth rate of the entire population. Note that ${\hat{p}}(e)$ that appears in the average fitness is the estimate of the true environmental probability, $p(e)$, obtained by averaging the likelihood across types.
This achieves a definition of fitness for each type $s$ as a variable probabilistic model of the environment, addressing the circularity problem. The picture of adaptation that emerges is explicitly relational and statistical: Selection relies on each phenotype having a generative model of the environment $p(e|s)$ and it is the goodness of this model, suitably averaged, that will translate into reproductive success.
@@
The fitness defined in Eq. 3 is a stochastic quantity because it is evaluated at (random) states of the environment. Consequently, we must discuss how to average it over the environment and what quantities are maximized by the population dynamics of Eq. 2. The connection between fitness, probability and information comes into focus when we consider that
@@
which is the instantaneous growth rate for type $s$. The first term shows the payoffs from the environment, the second measures environmental variability. The last term measures how close the $s$-type model of the environment is to the real thing, $p(e)$. Only the last term depends on the type and contributes to relative fitness.
When the population dynamics is iterated, the types with the highest average log-fitness are positively amplified. This emerges from the temporal dynamics because
$w_{s}(E)=\prod_{t=0}^{T}A(e_{t})p(e_{t}|s),\quad\rightarrow\ {\bar{\gamma}}_{s}=\sum_{i=1}^{T}\ln A(e_{t})p(e_{t}|s)\simeq\ln A+\sum_{e}p(e)\ln p(e)+\sum_{e}p(e)\ln\frac{p(e|s)}{p(e)}.$ (5)
Here, ${\bar{\gamma}}_{s}$ is the average growth rate of type $s$ over a long time $T$ and $\ln A$ is the average of $\ln A(e)$, see Supplementary Information. The long-term temporal average becomes the average over $e$ due to the law of large numbers. Thus, the population dynamics selects the types $s$ that maximize the average growth rate, which can be interpreted in terms of information quantities [50, 51]. The relative growth rate is maximized for $s$ such that the last term (which is the relative entropy and is negative definite) vanishes. This happens when the type’s fitness $p(e|s)\rightarrow p(e)$ becomes a perfect model of the real environment. The average of the type’s growth rates over the posterior $f(s|e)$, results in the mutual information $I(E,S)$, see Supplementary Information. In this sense, selection (1-3) converges on the type distribution that maximizes the information between population structure and the statistics of the environment.
Two additional consequences follow from the definition of fitness as a Bayesian likelihood. First, because fitness is a probability, statistically independent contributions combine as products. This means that fitness will only rarely be a linear combination of factors as is often assumed. Second, independent contributions do add up in the log-fitness, as they do in all information quantities. This makes the average log-fitness (or growth rate) the natural quantity for the analysis of complex evolutionary problems, even away from weak selection.
@@
In the problem of task switching, a population of agents is distributed to a set of discrete tasks, $s$, each associated with differential payoffs dependent on environmental states, $e$. Optimal evolutionary dynamics align these tasks with the environment to maximize rewards. For simplicity, we consider two tasks, denoted $s=0,1$, in a binary environment, $e=0,1$. The standard setup prescribes rewards, $Y_{s}(e)$, to each of these state combinations, Table 1A-B.
@@
The identification of population dynamics of this type with iterated Bayesian selection requires mapping the rewards $Y_{s}(e)$ to the likelihood function, $p(e|s)$. Because the likelihood is a conditional probability model, it is positive and normalized, meaning that any mapping to payoffs must satisfy such requirements. We want mappings that preserve the relative strength of the payoffs. This is achieved by writing the fitness, $w_{s}(e)=Ap(e|s)=Y_{s}(e)$. This requires that $Y_{s}(e)>0$ as assumed in Table 1 and that the factor of $A(e)$ is defined to ensure that none of the probabilities exceeds one and that the conditional probability is normalized. In general, the factor of $A$ is a function of the state of $e$, but since the example of Table 1 is symmetric, $A=5/2$ and $p(e=0|s=e)=p(e=1|s=e)=\frac{4}{5}$ and $p(e=0|s\neq e)=p(e=1|s\neq 0)=\frac{1}{5}$. We can write the fitness as
$w_{s}(e)=Y_{+}^{(1-s)(1-e)+se}Y_{-}^{(1-s)e+s(1-e)},\qquad\ln w_{s}=\ln Y_{+}+\ln\frac{Y_{+}}{Y_{-}}\left(2se-s-e\right).$ (6)
Note that we can now obtain the probability of $s$ that will maximize fitness over a long (stationary) history of the environment by taking expected values over $e$ and $s$. We define $f(s=1)=p$ and $p(e=1)=q$, and $f_{1|1}=f(s=1|e=1)$ and $f_{1|0}=f(s=1|e=0)$. The log-fitness is maximized when the quantity $(2f_{1|1}-1)q-p$ is maximal. This requires $f_{1|1}=1$, which implies $p=f_{1|1}q+f_{1|0}(1-q)=q$, which means that the population distribution of tasks is equal to the environmental probability.
We can also write the dynamics of selection by iterating the likelihood, given an environmental history and an initial distribution of traits $s$ in the population, see Supplementary information
@@
where $N_{e=s}$ is the number of times that $s=e$. For large times, the law of large numbers tells us that $N_{e=s}\rightarrow Tp(e,s)=Tp(s|e)p(e)$. The likelihood is maximized by taking $p(s=e|e)\rightarrow 1$, which means that $f(s)\rightarrow p(e)$, as derived above.
@@
Evolutionary game theory provides a general framework for the study of coevolutionary dynamics resulting from social interactions. The population now becomes an environment from which each agent derives its fitness. This contrasts with the task selection problem, where environmental dynamics are stationary and therefore independent of agent actions. The simplest situation is a mixture of two games: One game when the state of the environment $e=0$, and another when $e=1$. The total payoff over time will then be the result of agents recognizing which environmental situation they are in and adjusting their interactions accordingly. This also generalizes the task-switching problem to what we may call a game-switching situation.
We first demonstrate how the statistical treatment of fitness simplifies the analyses of dyadic games. Consider first the prisoners’ dilemma (PD), used for decades to analyze the evolution of cooperation [29, 20]. Table 1C-D shows the game’s payoff structure, where $s=0$ corresponds to defection and $s=1$ to cooperation. We consider two players $s_{1},s_{2}$ each with probability cooperation $p_{1},p_{2}$, respectively. Payoffs are chosen such that $W>R>L>P>0$. We write $s_{1}$’s fitness as
@@
Note that there is no unique prescription for the population dynamics of games [52]. Eq. 8 gives stochastic pay-offs at each interaction rather than on average, which would requires additional assumptions [52, 53]. This leads to
$\ln w_{s_{1}}=a-cs_{1}+bs_{2}+ds_{1}s_{2},\qquad a=\ln AL,\ c=\ln\frac{L}{P},\ b=\ln\frac{W}{L},\ d=\ln\frac{RL}{PW}.$ (9)
This form of log-fitness coincides with the typical parameterization of fitness as a linear function of costs and benefits of cooperation in the classical problem of altruism. Specifically, there is a fitness cost $c>0$ of cooperating when $s_{1}=1$, an independent benefit $b>0$ from the cooperation of others, $s_{2}=1$. There is also an extra benefit (if $d>0$) when both agents cooperate, which was considered by Queller [23, 22]. Note also that $d=c-\delta<c$, with $\delta=\ln W/R>0$. We see that, in a single-shot game, $s_{1}=0$ (defection) results in higher log-fitness than cooperation, even when the other player cooperates ($s_{2}=1$). This is the familiar Nash equilibrium, because the log-fitness of $s_{2}$ is identical with labels interchanged.
@@
By contrast, the goal of evolutionary game theory is the identification of selected strategies over long histories of interaction, which is analyzed via averaging Eq. 9. There are three distinct situations with growing levels of complexity. First, when the two agents act independently. Second, when they are part of a fully mixing population. And third, when their phenotypes become (statistically) dependent. Each of these general situations is associated with different population outcomes.
@@
The quantity in the square bracket is always positive, so $p_{1}=0$ maximizes the fitness independently of $p_{2}$. This means that the agent should always defect, as expected. Here, statistical independence means intuitively that the other agent does not react, even when it is exploited. In this sense, it must correspond to the single shot game and has the same optimal solution. Note also that, because of statistical independence, there can be no information between the agents.
@@
Now consider a seemingly small but important change in these assumptions. We consider $p_{1}=p_{2}=p$, which is frequently assumed in population genetics. It is dubbed random mating (in diploid populations) or a fully mixing population [1]. This results in
$\ln w_{s_{1}}=a-(c-b)p+dp^{2},$ (11)

### v2

@@ Abstract
Evolution is the process of optimal adaptation of biological populations to their living environments. This is expressed via the concept of fitness, defined as relative reproductive success. However, it has been pointed out that this definition is incomplete and logically circular. To address this issue, several authors have called for new ways to specify fitness explicitly in terms of the relationship between phenotypes and their environment. Here, we show that fitness, defined as the likelihood function that follows from mapping population dynamics to Bayesian learning, provides a general solution to this problem. We show how probabilistic models of fitness can easily be constructed in this way, and how their averages acquire meaning as information. We also show how this approach leads to powerful tools to analyze challenging problems of evolution in variable environments, game theory, and selection in group-structured populations. The approach is general and creates an explicit formal bridge between population dynamics, statistical learning theory, and emerging models of artificial intelligence.
@@ Introduction
Evolutionary modeling and theory have advanced rapidly over the past century incorporating the theory of replicators [1, 2, 3, 4], sources of intragenomic variation [5, 6], environmental mechanisms of genetic variation [7, 8], and socio-ecological interactions via game theory [9] and the Price equation[10].
@@
All of these approaches rely on the concept of fitness. Typically, fitness is defined as reproductive success, estimated by ratios of offspring to parents [1, 11]. Higher relative fitness of a type results in higher prevalence in the population over time. Most studies of evolutionary dynamics adopt this framework to identify better-adapted phenotypes in ecosystems, theoretical models [12, 13, 14], and laboratory studies of gene expression  [15].
However, it has been pointed out that this definition of fitness is incomplete and logically circular. This is because it does not explicitly refer to environmental adaptation [16, 17]. This omission leads to a number of interrelated problems.
@@
The first problem deals with the nature of adaptive capacity, which must be a persistent property of phenotypes in variable environments and cannot depend on single events [15, 18]. Any solution to this problem must explain how fitness emerges from the accumulation of exposures over time and environmental states. This issue also surfaces when modeling strategic interactions, such as in evolutionary game theory [19, 20].
@@
Related to temporal and population averaging, is the crucial question of how fitness combines across scales. It is well known in replicator models that contributions to fitness multiply [1, 11, 21]. This raises questions about approximations to fitness as a linear function of traits in both modeling and empirical analysis [22, 23]. The linearity of contributions to fitness is often justified in situations of weak selection. This is commonly assumed in the analysis of kin and group selection [22, 24] and in gene expression experiments [25]. Because many results and methods are based on the linearity of fitness, it becomes important to place this assumption on a stronger basis and develop methods that are suitable for circumstances when selection is not weak  [26, 27, 28, 29].
@@
Other problems arise from explicit environmental or social selection, for example, in the context of evolutionary game theory. An illustration is the difficulty of deriving definite predictions from strategic interaction games for the evolution of cooperation, or altruism [30, 24, 29]. A related problem is the proliferation of Hamilton-type rules for similar models [31]. These complications have led to the extensive use of computer simulations with different fitness definitions [32, 33], network and group structures [34, 35], and strategy spaces [36, 37, 38, 39]. This important body of work has emphasized repeated interactions [40, 41] and different types of reciprocity [31, 30] as fundamental conditions for cooperation to emerge and the need for group formation and intergroup competition [42, 22, 43]. However, these findings remain contested and relatively disorganized [24], in part because of the difficulty of analyzing the impact of assumptions and modeling choices.
@@
Here, we show that these conceptual and practical problems share a common root in the theoretical foundations of fitness. We demonstrate how they can be addressed by re-defining fitness as an optimal statistical adaptation to a probabilistic environment. Population dynamics under selection are then identified with iterated Bayesian updates of type frequencies. This identification has been recently proposed by several authors [44, 45, 46, 47], who noted its mathematical consistency and some of its conceptual implications. We take these ideas further by showing how they lead to the interpretation of fitness as a Bayesian likelihood, which is the optimal predictor of environmental states. We also show how fitness, hence constructed, allows us to build and analyze models of phenotype-environment selection in simple and clear ways for a variety of complex problems. This includes the effects of probabilistic environments in task selection, evolutionary dyadic games, and selection in group-structured populations.
@@
Phase spaces of stable strategies for these problems predict evolutionary transitions, expressed in the language of probability and information. Perhaps most satisfying, this approach shows how selection in evolutionary dynamics is a form of optimal learning theory in populations and how fitness maximization corresponds to the fundamental principle of maximizing information between population structure and the statistics of ecologically relevant environments.
@@ Results
Population replicators and growth dynamics
We define a structured population evolving under replicator dynamics. The population is made up of different types, $s$, such that the total population size, $N=\sum_{s}N_{s}$, where $N_{s}$ is the population of each type at a given time. The temporal evolution follows the standard replicator equation for a haploid population in discrete time $\tau$, $N_{s}(t)=w_{s}N_{s}(t-\tau),$ so that
$N(t)=\sum_{s}N_{s}(t)=\sum_{s}w_{s}N_{s}(t-\tau)=wN(t-\tau),$ (1)
where $w_{s}\geq 0$ is the type’s fitness. The type $s$ will grow if $w_{s}>1$ and decrease for $w_{s}<1$. This leads to the growth equation for the total population with average fitness $w=\sum_{s}w_{s}f_{s}$. Here, $f_{s}=\frac{N_{s}}{N}$ is the fraction of the population of type $s$. The total population will grow if $w>1$. The growth rate, $\gamma_{s}$, is related to fitness by exponentiation $w_{s}=e^{\gamma_{s}\tau}\simeq 1+\gamma_{s}\tau$, and, correspondingly $\gamma_{s}=\frac{1}{\tau}\ln w_{s}\simeq\frac{w_{s}-1}{\tau}$. These relationships apply in continuous time while the approximations are first-order series expansions. In the following, we set $\tau=1$ for simplicity.
@@
Fitness refers to optimal adaptation to an external environment, $e$. However, Eq. 1 and the standard set-up of population genetics [1] make no explicit reference to $e$: We now show how these perspectives can be reconciled.
@@
$f_{s}(t)=\frac{N_{s}(t)}{N(t)}=\frac{w_{s}}{w}\frac{N_{s}(t-1)}{N(t-1)}=\frac{w_{s}}{w}f_{s}(t-1).$ (2)
The mapping to Bayes’ theorem follows as
$f_{s}(t)=\frac{w_{s}}{w}f_{s}(t-1)\rightarrow p(H|D)=\frac{p(D|H)}{p(D)}p(H).$ (3)
Several authors have recently noted that the replicator equation under pure selection, Eq. 2, maps exactly to Bayes relation [47, 46], written on the right side of Eq. 3. Bayes theorem gives the exact update of an initial probability that a hypothesis $h\in H$ is true, $p(h)$ (the prior), given new evidence $D$. This is written as the updated (posterior) probability $p(H|D)$ over all $h\in H$. Because Bayes’ theorem is an identity that follows from the rules of probability, this rule is the optimal way to incorporate new evidence to update any initial probability [48]. As $p(H|D)$ is more constrained than $p(H)$ the data effectively selects specific hypotheses as more likely in a process where information is learned from the data about the hypothesis space. The key quantity in this procedure is $p(D|H)$, known as the likelihood. Although it looks like another probability, the likelihood evaluated over the data and the hypotheses is the engine of selection. It is interpreted in Bayesian estimation as a probabilistic model predicting the data given each hypothesis. The remaining factor is a normalization and does not introduce any new quantity because $p(D)=\sum_{h}p(D|h)p(h)$.
@@
This identification redefines fitness as an explicit model of environmental adaptation without additional assumptions. To see this, let us map the data, $D$, to environmental states, $D\rightarrow e$, and hypotheses to population types, $H\rightarrow s$. We then identify $f_{s}(t)=f(s|e)$, where $e$ is the state of the environment at time $t$, leading to
$f(s|e)=\frac{p(e|s)}{{\hat{p}}(e)}f(s)\rightarrow w_{s}(e)=A(e)p(e|s),\quad$ (4)
where the fitness can be expressed in terms of the likelihood,
$w(e)=\sum_{s}w_{s}(e)f(s)=A(e)\sum_{s}p(e|s)f(s)=A(e){\hat{p}}(e).$ (5)
This leads to the identification of fitness with the Bayesian likelihood, $w_{s}(e)=A(e)p(e|s)$, up to a multiplicative constant in the phenotypes, $A(e)$. The factor of $A(e)$, is necessary because fitness, defined in Eq. 1, $w_{s}=N_{s}(t)/N_{s}(t-1)>1$, when there is growth. The factor $A$ disappears in relative fitness, but contributes to absolute fitness and to the growth rate of the entire population. Note that ${\hat{p}}(e)$ in the average fitness is the estimate of the true environmental probability, $p(e)$, obtained by averaging the likelihoods across types.
This achieves a definition of fitness for each type $s$ as a variable probabilistic model of the environment addressing the circularity problem. The picture of adaptation that emerges is explicitly relational and statistical: Selection relies on diverse phenotypes as alternative generative models of the environment, $p(e|s)$. The goodness of these models versus the statistics of events in the environment over time results in relative reproductive success.
@@
The fitness defined in Eq. 5 is a stochastic quantity because it is evaluated at successive (random) states of the environment. Consequently, we must discuss how to average it and what quantities are maximized by the population dynamics of Eq. 2. The connection between fitness, probability and information comes into focus when we consider that
@@
which is the instantaneous growth rate, $\gamma_{s}$. The first term shows the payoffs from the environment, the second measures environmental variability, the third term measures how close the $s$-type model of the environment is to the real thing, $p(e)$, as we show next. Only this last term depends on the type and contributes to relative fitness.
When the population dynamics is iterated over a long time $T$, the types with the highest average log-fitness over such period are positively amplified. This emerges from the temporal dynamics over $T$ because
$w_{s}(E)=\prod_{t=0}^{T}A(e_{t})p(e_{t}|s)\rightarrow\ {\bar{\gamma}}_{s}=\frac{1}{T}\sum_{i=1}^{T}\ln A(e_{t})p(e_{t}|s).$ (7)
It follows that
${\bar{\gamma}}_{s}\simeq\ln A+\sum_{e}p(e)\ln p(e)+\sum_{e}p(e)\ln\frac{p(e|s)}{p(e)}.$ (8)
Here, ${\bar{\gamma}}_{s}$ is the average growth rate of type $s$ over a long temporal history. The approximation converges to the actual rate because the long-term temporal trajectory effectively averages over $e$ due to the law of large numbers. All three terms have interesting meanings. The quantity $\ln A$ is the average of $\ln A(e)$, see Supplementary Information. It is a property of the environment, setting the magnitude of the largest possible growth rate. The second term is recognizable as the negative Shannon entropy of the environment, $-H(E)=\sum_{e}p(e)\ln p(e)$: it is always negative and larger in more variable environments, which therefore entail a growth penalty because they are more difficult to predict. The last term is $-D(p(E)||p(E|s)\leq 0$, which is the negative relative entropy, or Kullback-Leibler divergence [50]. This quantity measures the distance in probability between the two distributions, meaning that it is maximized for the type $s$ with the perfect (or best) model of the environment $p(e|s)\rightarrow p(e)$. Other types will therefore incur a relatively smaller growth rate. Thus, the population dynamics selects the types $s$ that maximize the average growth rate, which can be interpreted in terms of information quantities [51, 52]. The average of the type’s growth rates over the posterior $f(s|e)$, results is the mutual information $I(E,S)$ between the population distribution over types and the probability of environmental states, see Supplementary Information. In this sense, selection (1-5) converges on the type distribution that maximizes the information between population structure and the statistics of the environment.
Two additional consequences follow from the definition of fitness as a Bayesian likelihood. First, because fitness is a probability, statistically independent contributions combine as products. This means that fitness will only rarely be a linear combination of factors, as often assumed. Second, independent contributions do add up in the log-fitness, as in all information quantities. This makes the average log-fitness (growth rate) the natural quantity for analyzing complex evolutionary problems, especially away from weak selection.
@@
In the problem of task switching, a population of agents is distributed to a set of discrete tasks, $s$, each associated with differential payoffs dependent on environmental states, $e$. Optimal evolutionary dynamics align these tasks with the environment to maximize rewards. For simplicity, we consider two tasks, denoted $s=0,1$, in a binary environment, $e=0,1$. The standard setup prescribes rewards, $Y_{s}(e)$, to each state combination, Table 1A-B.
@@
The identification of population dynamics of this type with iterated Bayesian selection requires mapping the rewards $Y_{s}(e)$ to the likelihood function, $p(e|s)$. Because the likelihood is a conditional probability, it is positive and normalized, meaning that any mapping to payoffs must satisfy such requirements. We want mappings that preserve the relative strength of the payoffs. This is achieved by writing the fitness, $w_{s}(e)=Ap(e|s)=Y_{s}(e)$. This requires that $Y_{s}(e)>0$ as assumed in Table 1 and that the factor of $A(e)$ is defined to ensure that none of the probabilities exceeds one and that the conditional probability is normalized. In general, the factor of $A$ is a function of the state of $e$, but since the example of Table 1 is symmetric, $A=5/2$ and $p(e=0|s=e)=p(e=1|s=e)=\frac{4}{5}$ and $p(e=0|s\neq e)=p(e=1|s\neq 0)=\frac{1}{5}$. We can write the fitness as
$w_{s}(e)=Y_{+}^{(1-s)(1-e)+se}Y_{-}^{(1-s)e+s(1-e)}.$ (9)
The logarithm of the fitness (log-fitness) is then
$\ln w_{s}=\ln Y_{+}+\ln\frac{Y_{+}}{Y_{-}}\left(2se-s-e\right).$ (10)
We can now obtain the probability of $s$ that maximizes fitness over a long (stationary) history of the environment by taking expected values over $e$ and $s$. We define $f(s=1)=p$ and $p(e=1)=q$, and $f_{1|1}=f(s=1|e=1)$ and $f_{1|0}=f(s=1|e=0)$. The log-fitness is maximized when the quantity $(2f_{1|1}-1)q-p$ is maximal. This requires $f_{1|1}=1$, which implies $p=f_{1|1}q+f_{1|0}(1-q)=q$. Hence, the population task distribution evolves to mirror the environmental probability.
We can also write the dynamics of selection by iterating the likelihood, given an environmental history and an initial distribution of traits $s$, see Supplementary Information
@@
where $N_{e=s}$ is the number of instances that $s=e$ over $T$ times. For long times, the law of large numbers tells us that $N_{e=s}\rightarrow Tp(e,s)=Tp(s|e)p(e)$. The likelihood is maximized by taking $p(s=e|e)\rightarrow 1$, which means that $f(s)\rightarrow p(e)$, as derived above.
@@
Evolutionary game theory provides a general framework for the study of coevolutionary dynamics resulting from social interactions. The population now becomes an environment from which each agent derives its fitness. This contrasts with the task selection problem, where environmental dynamics are stationary and therefore independent of the agent’s actions.
We first demonstrate how the statistical treatment of fitness simplifies the analysis of dyadic games. Consider the prisoners’ dilemma (PD), used for decades to analyze the evolution of cooperation [30, 19]. Table 1C-D shows the game’s payoff structure, where $s=0$ denotes defection and $s=1$ cooperation. We consider two players $s_{1},s_{2}$ each with cooperation probability $p_{1},p_{2}$, respectively. Payoffs are such that $W>R>L>P>0$. We write $s_{1}$’s fitness as
@@
Note that there is no unique prescription for the population dynamics of games [53]. Eq. 12 gives stochastic pay-offs at each interaction rather than on average, which would require additional assumptions [53, 54]. This leads to
$\ln w_{s_{1}}=a-cs_{1}+bs_{2}+ds_{1}s_{2},$ (13)
where $a=\ln AL$, $c=\ln\frac{L}{P}$, $b=\ln\frac{W}{L}$, and $d=\ln\frac{RL}{PW}$. This polynomial log-fitness coincides with the typical parameterization of fitness as a linear function of costs and benefits of cooperation in the classical problem of altruism [30, 19]. Specifically, there is a fitness cost, $c>0$, of cooperating, and an independent benefit $b>0$ from the cooperation of others. There is also an extra benefit (if $d>0$) when both agents cooperate, which was considered by Queller [23, 22]. Note also that $d=c-\delta<c$, with $\delta=\ln W/R>0$. In a single-shot game, $s_{1}=0$ (defection) results in higher log-fitness even when the other player cooperates ($s_{2}=1$). This is the game’s familiar Nash equilibrium, because the log-fitness of $s_{2}$ is identical with labels interchanged.
@@
By contrast, the goal of evolutionary game theory is the identification of selected strategies over long histories, which is analyzed via averaging Eq. 13. There are three distinct situations with growing levels of complexity. First, when the two agents act independently. Second, when they are part of a fully mixing population. And third, when their phenotypes become (statistically) dependent. Each of these general situations is associated with different outcomes.
@@
The quantity in the square bracket is always positive, so $p_{1}=0$ maximizes the fitness independently of $p_{2}$. This means that the agent should always defect, as expected. Here, statistical independence means intuitively that the other agent does not react, even when it is exploited. In this sense, it must correspond to the single shot game and has the same optimal solution. Note also that statistical independence means that there cannot be any information between the agents.
@@
Now consider a seemingly small but important change in these assumptions. We consider $p_{1}=p_{2}=p$, which is frequently assumed in population genetics. This is often dubbed random mating (in diploid populations) or a fully mixing population [1], resulting in
$\ln w_{s_{1}}=a-(c-b)p+dp^{2}.$ (15)